A gun is fired with a velocity of 150m|s at an angle of 30°to the horizontal calculate its time of flight

1
Expert's answer
2022-03-11T14:02:56-0500

The time of flight is given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"t = \dfrac{2u\sin\theta}{g}"

where "u = 150m\/s" is the initial speed, "\theta = 30\degree" is the launch angle, "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"t = \dfrac{2\cdot 150m/s\cdot \sin30\degree}{9.8m/s^2} \approx 15s"

Answer. 15s.